# A note on generalized tensor CUR approximation for tensor pairs and tensor triplets based on the tubal product

### v2

@@ 12       $\widehat{\underline{\mathbf{S}}}_{i}=\widehat{\underline{\mathbf{S}}}_{I_{3}-i+2}$;
3       ${\bf p}(j)={\arg\max}_{1\leq i\leq I_{2}}\|\underline{\bf Z}(i,j,:)\|_{2}$
4       ${\bf s}_{1}(j)=\arg\max_{1\leq i\leq I_{1}}\|\underline{\bf U}(i,j,:)\|_{2}$
5       ${\bf s}_{2}(j)=\arg\max_{1\leq i\leq I_{4}}\|\underline{\bf V}(i,j,:)\|_{2}$
@@ 6 ${\underline{\mathbf{G}}}_{4}=\widehat{\underline{\mathbf{Y}}}({\bf s}_{2},:,:)$;
In what follows and similarly to [2], we show how the GTCUR of the tensor pairs $(\underline{\bf X},\underline{\bf Y})$ is related to the TCUR of the tensors $\underline{\bf X}*\underline{\bf Y}^{-1}$ and $\underline{\bf X}*\underline{\bf Y}^{{\dagger}}$ for invertible or full tubal rank tensor $\underline{\bf Y}$.
@@
    b) As a special case, if $\underline{\bf Y}=\underline{\bf I}$, then the GTCUR of the tensor pairs $(\underline{\bf X},\underline{\bf Y})$ is the same as the TCUR of the data tensor $\underline{\bf X}$. We mean that (10) is the same as (13).
@@ 6 Tensor CUR approximation for tensor triplets based on the t-product
Figure 3: Visualization of the generalized tensor CUR (GTCUR) approximation for the tensor triples $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$. The indices of the selected horizontal slices of the tensors $\underline{\bf X}$ and $\underline{\bf Y}$ are identical, but but this is not necessarily true for the lateral slices. Additionally, the indices of the selected lateral slices of $\underline{\bf X}$ and $\underline{\bf Z}$ are the same, however, this is not always the case for the sampled horizontal slices.
@@
where $\underline{\bf S}\in\mathbb{R}^{I_{1}\times R\times I_{3}},\,\widehat{\underline{\bf S}}\in\mathbb{R}^{I_{1}\times R\times I_{3}},\,\underline{\bf P}\in\mathbb{R}^{I_{2}\times R\times I_{3}}$ and $\widehat{\underline{\bf P}}\in\mathbb{R}^{I_{4}\times R\times I_{3}},\,R\leq\min(I_{1},I_{2},I_{4},I_{5})$ are the index selection tensors yielded by sampling some lateral/horizontal slices of the identity tensor according to the sampled lateral/horizontal slices from the data tensors $\underline{\bf X},\underline{\bf Y},$ and $\underline{\bf Z}$. Alternative representations for (19)-(21) are
$\underline{\bf X}$$\approx$ $\underline{\bf X}(:,{\bf p},:)*\underline{{\bf U}}_{1}*\underline{\bf X}({\bf s},:,:),$ (22)
$\underline{\bf Y}$$\approx$ $\underline{\bf Y}(:,{\bf p}_{1},:)*\underline{{\bf U}}_{2}*\underline{\bf Y}({\bf s},:,:),$ (23)
$\underline{\bf Z}$$\approx$ $\underline{\bf Z}(:,{\bf p},:)*\underline{{\bf U}}_{3}*\underline{\bf Z}({\bf s}_{1},:,:),$ (24)
@@
The intuition behind the GTCUR for tensor triplets also comes from the tensor-based restricted t-SVD (t-RSVD) which is a direct generalization of the classical RSVD [7] and can be stated as follows. Let $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, $\underline{\bf Y}\in\mathbb{R}^{I_{1}\times I_{4}\times I_{3}}$, and $\underline{\bf Z}\in\mathbb{R}^{I_{5}\times I_{2}\times I_{3}}$ be given. Generalizing the RSVD to tensors based on the t-product guarantees the following decompositions
@@
Algorithm 8 provides a summary of the process for computing the GTCUR for tensor triples. Lines 6-8 can be efficiently computed in the Fourier domain and similar algorithms like Algorithm 6 can be developed for this computation. The t-RSVD of the tensor triplets $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$ provides the common tensor factors $\underline{\bf L},\,\underline{\bf W}$ and extra tensor factors $\underline{\bf U}$ and $\underline{\bf V}$. These bases can be used to sample lateral and horizontal slices of the tensor triplets $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$. Since we have the common tensor factors $\underline{\bf L}$ and $\underline{\bf W}$, the indices, which are used to sample the horizontal slices of the data tensors $\underline{\bf X}$ and $\underline{\bf Z}$ are the same, while the indices for the selection of lateral slices of the data tensors $\underline{\bf X}$ and $\underline{\bf Y}$ are identical. Nevertheless, the tensor bases $\underline{\bf U},\,\underline{\bf V}$ are used to sample lateral slice indices of the tensor $\underline{\bf Y}$ and horizontal slice indices of the tensor $\underline{\bf Z}$. The visualization of this approach is described in Figure 3. So, the idea is to compute the t-RSVD of the tensor triplets $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$ and to find the corresponding tensor factors $\underline{\bf U},\,\underline{\bf V},\,\underline{\bf Z},\,\underline{\bf W}$ and use them to sample horizontal and lateral slice indices using the TDEIM algorithm. We should remark that the computation of the t-RSVD is demanding for large-scale data tensors. However he double GSVD, which is necessary to compute the t-RSVD of the tensor triplets, may be done quickly using the fast randomized GSVD and randomized GTSVD techniques provided in [37, 35]. The connection between the GTCUR for tensor triplets with the GTCUR for tensor pairs and the TCUR is described in the next theorem.
@@
    a) If $\underline{\bf Y}$ and $\underline{\bf Z}$ are nonsingular tensors, then the selected lateral and horizontal slice indices from the TCUR decomposition of $\underline{\bf Y}^{-1}*\underline{\bf X}*\underline{\bf Z}^{-1}$ are identical to the index vectors ${\bf p}_{1}$ and ${\bf s}_{1}$, respectively, obtained from a GTCUR approximation of the tensor triples $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$.
@@
    b) In the particular case where $\underline{\bf Y}=\underline{{\bf I}}$ and $\underline{\bf Z}=\underline{{\bf I}}$, the GTCUR approximation of $(\underline{\bf X},\underline{\bf I},\underline{\bf I})$ coincides with the TCUR decomposition of $\underline{\bf X}$. We mean that (10) is the same as (22).
@@
    c) For a special choice of $\underline{\bf Y}=\underline{\bf I}$, a GTCUR approximation of $(\underline{\bf X},\underline{\bf I},\underline{\bf Z})$ coincides with the GTCUR approximation of $(\underline{\bf X},\underline{\bf Z})$. Similar results can be stated for $\underline{\bf Z}=\underline{\bf I}$.
@@
a) From 25, it is not difficult to see that $\underline{\bf Y}^{-1}=\underline{\bf U}*{\underline{\bf D}}_{2}^{-1}*\underline{\bf L}^{-1},\,\underline{\bf Z}^{-1}=\underline{\bf W}*\underline{\bf D}_{3}^{-1}*\underline{\bf V}^{T}$, so we have
@@
Clearly, (27) and (28) shows that the GTCUR of the tensor pairs $(\underline{\bf X},\underline{\bf Z})$ provides the same indices for horizontal and lateral slice sampling as the GTCUR approximation of the tensor pairs $(\underline{\bf X},\underline{\bf I},\underline{\bf Z})$. So, the proof is completed. ∎
@@ 7 Conclusion
In this note, we showed how the tensor CUR (TCUR) approximation can be extended to tensor pairs and tensor triplets. We used the tensor Discrete Interpolatory Empirical Method (TDEIM) to generalize the TCUR to tensor pairs and tensor triplets. This extension is called the generalized TCUR (GTCUR) approximation. We established connections between some special cases of the GTCUR and the classical TCUR approximation. Efficient algorithms are presented to compute the GTCUR approximation for both tensor pairs and tensor triplets. We are investigating the theoretical and numerical aspects of the presented GTCUR approaches for tensor pairs and tensor triplets in a practical setting.
@@ 8 Conflict of Interest Statement
9 Data availability statement
This paper does not have any digital data available.

### v3

@@ 12       $\widehat{\underline{\mathbf{S}}}_{i}=\widehat{\underline{\mathbf{S}}}_{I_{3}-i+2}$;
3       ${\bf p}(j)={\arg\max}_{1\leq i\leq I_{2}}\|\underline{\bf Z}(i,j,:)\|$
4       ${\bf s}_{1}(j)=\arg\max_{1\leq i\leq I_{1}}\|\underline{\bf U}(i,j,:)\|$
5       ${\bf s}_{2}(j)=\arg\max_{1\leq i\leq I_{4}}\|\underline{\bf V}(i,j,:)\|$
@@ 6 ${\underline{\mathbf{G}}}_{4}=\widehat{\underline{\mathbf{Y}}}({\bf s}_{2},:,:)$;
In what follows and similar to [2], we show the connection between the GTCUR of the tensor pairs $(\underline{\bf X},\underline{\bf Y})$ and TCUR of the tensors $\underline{\bf X}*\underline{\bf Y}^{-1}$ and $\underline{\bf X}*\underline{\bf Y}^{{\dagger}}$ for non-invertible tensor, or of full rank $\underline{\bf Y}$ in Theorem 3.
@@
    b) As a special case, if $\underline{\bf Y}=\underline{\bf I}$, then the GTCUR of the tensor pairs $(\underline{\bf X},\underline{\bf Y})$ is the same as the TCUR of the data tensor $\underline{\bf X}$.
@@ 6 Tensor CUR approximation for tensor triplets based on the t-product
Figure 2: Visualization of the generalized tensor CUR (GTCUR) approximation for tensor triples $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$. The indices of the selected horizontal slices of the tensors $\underline{\bf X}$ and $\underline{\bf Y}$ are identical, but not those for the lateral slices. Also, the indices of the selected lateral slices of $\underline{\bf X}$ and $\underline{\bf Z}$ are the same, but not indices of the selected horizontal slices.
@@
where $\underline{\bf S}\in\mathbb{R}^{I_{1}\times R\times I_{3}},\,\widehat{\underline{\bf S}}\in\mathbb{R}^{I_{1}\times R\times I_{3}},\,\underline{\bf P}\in\mathbb{R}^{I_{2}\times R\times I_{3}}$ and $\widehat{\underline{\bf P}}\in\mathbb{R}^{I_{4}\times R\times I_{3}},\,R\leq\min(I_{1},I_{2},I_{4},I_{5})$ are the index selection tensors yielded by sampling some lateral/horizontal slices of the identity tensor according to the sampled lateral/horizontal slices, which are sampled from the data tensors $\underline{\bf X},\underline{\bf Y},$ and $\underline{\bf Z}$. Alternative representations for (18)-(20) are
$\underline{\bf Y}$$\approx$ $\underline{\bf Y}(:,{\bf p}_{1},:)*\underline{{\bf U}}_{2}*\underline{\bf Y}({\bf s},:,:),$
$\underline{\bf Z}$$\approx$ $\underline{\bf Z}(:,{\bf p},:)*\underline{{\bf U}}_{3}*\underline{\bf Z}({\bf s}_{1},:,:),$
@@
The intuition behind the GTCUR for tensor triplets also comes from the tensor-based restricted t-SVD (t-RSVD) which is a direct generalization of the classical RSVD [7] to tensors and can be stated as follows. Let $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, $\underline{\bf Y}\in\mathbb{R}^{I_{1}\times I_{4}\times I_{3}}$, and $\underline{\bf Z}\in\mathbb{R}^{I_{5}\times I_{2}\times I_{3}}$ be given. Generalizing the RSVD to tensors based on the t-product guarantees the following decompositions
@@
The procedure of the computation of the GTCUR for tensor triples is summarized in Algorithm 8. Lines 6-8 can be efficiently computed in the Fourier domain and similar algorithms like Algorithm 6 can be developed for this computation. The t-RSVD of the tensor triplets $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$ provides the common tensor factors $\underline{\bf L},\,\underline{\bf W}$ and extra tensor factors $\underline{\bf U}$ and $\underline{\bf V}$. These bases can be used to sample lateral and horizontal slices of the tensor triplets $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$. Since we have the common tensor factors $\underline{\bf L}$ and $\underline{\bf W}$, the indices, which are used to sample the horizontal slices of the data tensors $\underline{\bf X}$ and $\underline{\bf Z}$ are the same, while the indices for the selection of lateral slices of the data tensors $\underline{\bf X}$ and $\underline{\bf Y}$ are identical. Nevertheless, the tensor bases $\underline{\bf U},\,\underline{\bf V}$ are used to sample lateral slice indices of the tensor $\underline{\bf Y}$ and horizontal slice indices of the tensor $\underline{\bf Z}$. The visualization of this approach is described in Figure 2. So, the idea is to compute the t-RSVD of the tensor triplets $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$ and to find the corresponding tensor factors $\underline{\bf U},\,\underline{\bf V},\,\underline{\bf Z},\,\underline{\bf W}$ and use them to find the indices for selecting horizontal and lateral slices using the TDEIM algorithm. We should remark that the computation of the t-RSVD is demanding for the case of large-scale data tensors and the fast randomized GSVD algorithm proposed in [40] can be used for the computation of the double GSVD, which are required to compute the t-RSVD of the tensor triplets. The connection between the GTCUR for tensor triplets with the GTCUR for tensor pairs and the TCUR is described in the next theorem.
@@
    a) If $\underline{\bf Y}$ and $\underline{\bf Z}$ are nonsingular tensors, then the selected lateral and horizontal slice indices from the TCUR decomposition of $\underline{\bf Y}^{-1}*\underline{\bf X}*\underline{\bf Z}^{-1}$ are identical to the index vectors ${\bf p}_{1}$ and ${\bf s}_{1}$, respectively, obtained from a GTCUR decomposition of the tensor triples $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$.
@@
    b) In the particular case where $\underline{\bf Y}=\underline{{\bf I}}$ and $\underline{\bf Z}=\underline{{\bf I}}$, the GTCUR decomposition of $(\underline{\bf X},\underline{\bf I},\underline{\bf I})$ coincides with the TCUR decomposition of $\underline{\bf X}$.
@@
    c) For a special choice of $\underline{\bf Y}=\underline{\bf I}$, a GTCUR approximation of $(\underline{\bf X},\underline{\bf I},\underline{\bf Z})$ coincides with the GTCUR decomposition of $(\underline{\bf X},\underline{\bf Z})$. Similar results can be stated for $\underline{\bf Z}=\underline{\bf I}$.
@@
a) From 21, it is not difficult to see that $\underline{\bf Y}^{-1}=\underline{\bf U}*{\underline{\bf D}}_{2}^{-1}*\underline{\bf L}^{-1},\,\underline{\bf Z}^{-1}=\underline{\bf W}*\underline{\bf D}_{3}^{-1}*\underline{\bf V}^{T}$, so we have
@@
Clearly, (23) and 24 shows that the GTCUR of the tensor pairs $(\underline{\bf X},\underline{\bf Z})$ provides the same indices for horizontal and lateral slice sampling as the GTCUR decomposition of the tensor pairs $(\underline{\bf X},\underline{\bf I},\underline{\bf Z})$. So, the proof is completed. ∎
@@ 7 Conclusion
In this note, we showed how the tensor CUR (TCUR) approximation can be extended to tensor pairs and tensor triplets. We use the tensor Discrete Interpolatory Empirical Method (TDEIM) to generalize the TCUR to tensor pairs and tensor triplets. This extension is called the generalized TCUR (GTCUR) method. We established connections between some special cases of the GTCUR and the classical TCUR approximation. Efficient algorithms are presented to compute the GTCUR approximation for both tensor pairs and tensor triplets. We are investigating the theoretical and numerical aspects of the presented GTCUR approaches for tensor pairs and tensor triplets in a practical setting.